I keep confusing myself I do not understand the slope formula and I have an assignment due today. The equation of a line perpendicular to the given line but passing through the given point y=-2/3x+2 (9,-3)
Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

The equation of a line perpendicular to the given line but passing through the given point y=-2/3x+2 (9,-3)
-------------------------
That line has a slope of -2/3
----
Your line must have a slope of 3/2 because their product must be -1.
-----------------------
Form of your equation: y = mx + b
----
Solve for "b":
-3 = (3/2)(9) + b
--------
-6/2 = 27/2 + b
----
b = -33/2
-------
Equation:
y = (3/2)x - (33/2)
======================
Cheers,
Stan H.
======================

Answer by richwmiller(17219)

y=mx+b
m is the slope and
b is the value of y when x=0 It is often called the y intercept
Parallel lines have the same slope
the product of the slopes of perpendicular lines is -1
so if the slope of one line is 4 the slope of the other is -1/4 (4*-1/4=-1)
The slope of the line y=-2/3x+2 is -2/3
So the slope of the line perpendicular line is 3/2
3/2*(-2/3)=-1
point (9,-3)
y=mx+b
m = 3/2, x = 9, y = -3
-3=3/2*(9)+b
b = -33/2,
y=3x/2-33/2
the two perpendicular lines cross at x = 111/13, y = -48/13